A racing car passes through the horizontal turn at a speed of 180 km / h. Calculate the mass of the car, when it is known that the centripetal force at the sliding threshold is 15000 N, at the turning radius 200 m. Determine the coefficient of sliding friction.


1
Expert's answer
2021-01-30T06:25:17-0500

"180(km\/h)=50(m\/s)"


"F_f=m\frac{v^2}{R}\to m=\frac{F_f\cdot R}{v^2}=\frac{15000\cdot 200}{50^2}=1200\ (kg)" . Answer


"F_f=\mu N=\mu mg\to\mu=\frac{F_f}{mg}=\frac{15000}{1200\cdot9.8}=1.275" . Answer